# Whole-body end-effector pose tracking

### v1

@@ I INTRODUCTION
Over the past decade, recent algorithmic advancements have substantially increased legged robots’ ability to traverse complex, cluttered environments and human-designed infrastructures, such as stairs and slopes [1, 2, 3, 4]. Despite these advancements, their practical applicability remains constrained by their limited manipulation capabilities. Most field operations with legged robots involve minimal environmental interactions through visual inspections and passive load transportation. Thus, combining a legged robot’s mobility with the ability to perform manipulation tasks is critical for enhancing their applications to more real-world scenarios.
Compared to fixed base counterparts, integrating an arm onto a legged mobile platform significantly complicates the controller design due to increased degrees of freedom, redundancy and highly non-linear dynamics. To address this, the research community has mainly explored model-based and learning-based control strategies.
@@
More recently, Reinforcement Learning (RL) has emerged as a robust alternative, directly learning control policies through interactions in simulation, for locomotion [9, 1, 10] and whole-body control [11]. Unlike MPC’s two-stage planning and tracking, this approach creates end-to-end control policies that map sensory inputs directly to motor commands, enabling improved resilience to external disturbances due to the introduction of environmental variability during training [12, 13, 14]. Research on legged robots, whether using legs [15, 16] or attached arms [11, 17], demonstrate that RL techniques can achieve effective end-effector position tracking across a large workspace through agile whole-body behavior. Although effective in outdoor and slippery terrains, these approaches are unsuitable beyond basic flat terrain and lack orientation tracking.
Figure 2: Our whole-body controller has precise end-effector pose tracking in a large workspace on flat terrain and stairs.
Figure 3: Training setup: The training process begins with data collection, where we gather (A) the terrain mesh and compute a coarse terrain height map representing the terrain surface, (B) pre-sample 10000 end-effector pose commands with a fixed base, and (C) collect base poses and joint angles using a pre-trained locomotion policy to initialize robots. During training, a command is sampled from (B), slightly transformed, and checked for collisions with the terrain. If collision-free, it is concatenated with observations and input to the policy; otherwise, a new command is sampled. The policy is trained in simulation with 4000 robots in parallel, outputting joint actions.
In this work, we present a general-purpose whole-body controller for legged robots with an attached arm, designed to provide stable end-effector pose tracking across a large operational space. Our approach includes a terrain-aware sampling strategy for end-effector pose commands, and for the robot’s initial configuration to ensure a smooth transition from a locomotion policy to the proposed whole-body controller. Experimental results show that the controller achieves precise tracking across varying terrains, such as stairs.
Our key contributions are as follows:
  • An RL-based whole-body controller for 6-DoF end-effector pose tracking for quadrupeds with an arm.
  • We showcase the learned controller’s tracking capabilities over challenging terrains and its robustness to external disturbances.
  • Compared to model-based controllers and existing RL works, we show that our learned controller achieves a higher tracking accuracy and enlarged pose reachability both in simulation and on real hardware, reaching a pose tracking accuracy of $2.64~{}$\mathrm{cm}$$ and $3.64\degree$.
@@ II-A Model-based whole-body control
Formulating the whole-body planning and control of legged mobile manipulators into a single optimization problem avoids treating the arm as an external disturbance to the base [18, 19] and helps ensure tight coordination between the base and the arm [20, 21]. Belliscoso et al.  [5] propose an online ZMP-based motion planning framework that relies on an inverse dynamics model to track generated operational-space references. In contrast, Sleiman et al.  [6] provide an unified MPC formulation for whole-body dynamic locomotion and manipulation planning. Chiu et al.  [22] further extend this formulation to account for self and environment collisions during the receding-horizon control. To increase the robustness towards external disturbances proactively, Ferrolho et al.  [23, 24] incorporate a robustness metric into the trajectory optimization to compute offline plans for interactions under worst-case scenario. While these approaches are effective for end-effector control of legged mobile manipulators, they rely on an available system model and pre-specified gait schedule. These assumptions make model-based approaches vulnerable to large unmodeled disturbances (for instance, a heavy grasped object), slippages, and unknown terrain models.
@@ II-B Learning-based whole-body control
More recently, RL has shown to be a powerful alternative to model-based approaches for legged robots, providing robust control policies for locomotion on unstructured terrains [1, 10] and whole-body manipulation [11, 15, 25]. In the context of task-space tracking of a legged mobile manipulator’s arm, existing works formulate the tracking problem in different ways, using different representations for the end-effector commands.
@@
Fu et al.  [11] train a combined RL policy for locomotion and manipulation. For the arm, they sample position commands in spherical coordinates and orientation commands uniformly in $\mathbb{SO}(3)$. However, their results are shown only for 3D position commands and on flat terrain. A follow-up comparison in [29] reports a poor pose tracking accuracy with this work as $22.2$\mathrm{cm}$$ and $66.22\degree$. Considering the importance of regulating forces during interactions, Portela et al.  [17] train an RL controller for tracking end-effector 3D position and 3D force simultaneously. However, they also demonstrate their work only on flat terrains. More recently, Ha et al.  [29] use a 3D Cartesian and 6D rotation representation [30] for the end-effector pose and train a low-level RL policy to track a trajectory of three-pose commands. While they achieve a position and orientation error of $2.12$\mathrm{cm}$$ and $3.35\degree$, respectively, the learned controller is mainly demonstrated over reference end-effector motions in front of the robot, which does not necessitate active base (leg) usage.
@@ III METHOD
We train a policy to track end-effector target poses with minimal foot displacement, intended for use alongside a locomotion policy. Figure 3 illustrates the overall training process, starting with a data collection step, gathering terrain information and pre-sampling end-effector poses. During training, pose commands are sampled, checked for collisions, and used to train the policy.
We use Isaac Lab [31] as a simulation environment to train our policy, and we deployed our controller on ALMA [5], which integrates the Anymal D robot from ANYbotics [32] with the Dynaarm from Duatic [33]. The robot has an inertial measurement unit providing body orientation and 18 actuators with joint position encoders. The main control loop and the state estimator are executed at 400 Hz, while our policy runs at 50Hz on an onboard computer. While we show results on ALMA, our method remains applicable across different robot embodiments.
@@ III-B Command sampling
The distribution of the pre-sampled commands is not uniform. We fit five z-aligned concentric cylinders, with the first cylinder solid and the subsequent cylinders having increasing radii with hollow sections. We determine their radii by first calculating the maximum radius in the $x$-$y$ plane from the 3D positions of all pre-sampled poses and dividing it into five equal parts. The distribution of poses across these cylinders, starting from the inner one, are as follows: 51%, 21%, 13%, 6%, and 2%. To enforce more spatially uniform sampling, we randomly select one of the bins and then sample a pose from within that selected bin.
@@ III-D Action and Observation Space
The observation, $o^{t}$, is augmented with the end-effector pose command for the policy input. This command is defined as the positional difference between the current and the target keypoints of the end-effector in the base frame $\prescript{b}{}{\!p_{ee}^{\text{cmd}}}\in\mathbb{R}^{9}$, where each of the three keypoints provides a 3D position vector, resulting in a 9-dimensional vector in total.
@@ III-F Terrains and Curriculum training
The robots are trained in simulation on four procedurally generated terrains: flat, randomly rough, discrete obstacles, and stairs, as defined in [35]. Gradually increasing task difficulty during training has been shown to enhance learning in previous works [1, 38, 35]. We employ a terrain curriculum similar to the one proposed in [35], but adapted for end-effector pose tracking. If the average position tracking error across the three commands within an episode is under $20$ $\mathrm{cm}$ when the task reward is active (during the final 2 seconds of each 4-second command), and the average orientation tracking error is under $20\degree$, the terrain difficulty increases after the next reset. The terrain difficulty decreases if the error exceeds $80$ $\mathrm{cm}$ and $120\degree$. Robots achieving the highest level are assigned to a random level to prevent catastrophic forgetting.
@@ IV-A2 Comparison to model-based control
We compare the tracking performance of our whole-body RL policy with the model-based whole-body MPC controller from prior work [22]. For this comparison, we use the same weight parameters as those in the original paper for the MPC controller, which has been optimized for the same robot. We first evaluate both controllers on flat terrain using the same 15 end-effector pose commands. For easy poses within the initial workspace (Figure 4A), both controllers perform similarly, with average errors of $1.22$ $\mathrm{cm}$ and $1.73\degree$ for the RL controller, and $1.52$ $\mathrm{cm}$ and $1.02\degree$ for the MPC controller. However, for poses in the expanded workspace (Figure 4B), the RL policy outperforms the MPC controller, showing average position errors of $0.96$ $\mathrm{cm}$ and $1.29\degree$ versus $2.71$ $\mathrm{cm}$ and $9.73\degree$ for the MPC controller. Figure 6 illustrates their tracking performance, highlighting the MPC controller’s limited whole-body behavior. While the MPC controller is primarily designed for flat terrains, we include one experiment on stairs with 15 poses under two stable base orientations. However, 30% of the tested orientations were unstable for the MPC controller and were excluded from this evaluation. In this experiment, the MPC controller reaches an average position error of $24.3$ $\mathrm{cm}$ and $45.47\degree$, whereas our RL policy achieves significantly better tracking accuracy with $3.5$ $\mathrm{cm}$ and $2.67\degree$. Furthermore, for the MPC controller, we observed that the arm often gets stuck near the base when moving between distant poses due to conflicting objectives of pose tracking and self-collision avoidance, a problem that our RL controller effectively avoids.
Figure 6: MPC controller end-effector tracking (left) shows $7.21$ cm position error and $21.91$ deg orientation error. Our whole-body RL controller (right) achieves $0.83$ cm cm and $0.61$ deg errors, thanks to its effective whole-body behavior. The command is indicated by the red arrow.
@@ V CONCLUSION
We have presented a whole-body RL-based controller for a quadruped with an arm that can reach even the most difficult poses. Our controller achieves high accuracy also on rough terrain (e.g., on stairs), which we show in real-world experiments with ANYmal with an arm. Additionally, our contributions include providing a formulation for learning pose tracking that is superior to existing methods with poor accuracy or only considering position tracking. Our hardware experiments show an average tracking accuracy of $2.64~{}$\mathrm{cm}$$ for position and $3.64\degree$ for orientation on challenging terrain. Future work includes integrating 3D representations of the environment into the pipeline to enable self-learned collision avoidance, as done for locomotion in [40].

### v2

@@ I INTRODUCTION
Over the past decade, algorithmic advancements have substantially increased legged robots’ ability to traverse complex, cluttered environments and human-designed infrastructures, such as stairs and slopes [1, 2, 3]. Despite these improvements, their practical applicability remains constrained by their limited manipulation capabilities. Most field operations with legged robots involve minimal environmental interactions, such as visual inspections and passive load transportation. Thus, combining a legged robot’s mobility with the ability to perform manipulation tasks is critical for enhancing their applications to more real-world scenarios.
Compared to fixed base counterparts, integrating an arm onto a legged mobile platform significantly complicates the controller design because of increased degrees of freedom, redundancy and highly non-linear dynamics. To address this, the research community has mainly explored model-based and learning-based control strategies.
@@
In contrast, the learning-based control strategy of Reinforcement Learning (RL) has emerged as a robust alternative, directly learning control policies through interactions in simulation for locomotion [8, 1, 3] and whole-body control [9]. RL enables improved resilience to external disturbances compared to MPC because of environmental variability during training [10, 11, 1]. Research on legged robots, whether using legs [12, 13] or attached arms [9, 14], demonstrate that RL techniques can achieve effective end-effector position tracking across a large workspace through agile whole-body behavior. Although effective in outdoor and slippery terrains, these approaches remain unsuitable beyond flat terrain and lack orientation tracking.
Figure 2: Our whole-body controller demonstrates precise end-effector pose tracking across a variety of challenging terrains, including soft mattresses, stairs and uneven natural ground.
Figure 3: The training process begins with data collection, where we gather (A) the terrain mesh and a coarse terrain height map, (B) 10000 pre-sampled end-effector pose commands with a fixed base, and (C) base poses and joint angles from a pre-trained locomotion policy to initialize robots. During training, a command from (B) is slightly transformed and checked for collisions with the terrain. If collision-free, it is concatenated with observations and input to the policy; otherwise, a new command is sampled. The policy is trained in simulation with 4000 robots in parallel, outputting joint actions.
To address these shortcomings, we propose a general-purpose whole-body controller for legged robots with an attached arm. The controller is designed to provide stable end-effector pose tracking across a large operational space. Our approach includes a terrain-aware sampling strategy for end-effector pose commands, and for the robot’s initial configuration to ensure a smooth transition from a locomotion policy to the proposed whole-body controller. Experimental results show that the controller achieves precise tracking across varying terrains, such as stairs and slopes. Our key contributions are as follows:
  • We propose an RL whole-body controller for 6-DoF end-effector pose tracking for quadrupeds with an arm.
  • We showcase the learned controller’s tracking capabilities over challenging terrains and its robustness when faced with external disturbances.
  • We compare our learned controller to model-based controllers and existing RL approaches, showing higher tracking accuracy and enlarged pose reachability, both in simulation and on hardware, reaching a pose tracking accuracy of $2.64~{}$\mathrm{cm}$$ and $3.64\degree$.
@@ II-A Model-based whole-body control
Formulating the whole-body planning and control of legged mobile manipulators into a single optimization problem avoids treating the arm as an external disturbance to the base [15, 16] and helps ensure tight coordination between the base and the arm [17, 18]. Belliscoso et al.  [4] propose an online ZMP-based motion planning framework that relies on an inverse dynamics model to track generated operational-space references. In contrast, Sleiman et al.  [5] provide a unified MPC formulation for whole-body dynamic locomotion and manipulation planning. Chiu et al.  [19] further extend this formulation to account for self and environment collisions during the receding-horizon control. To increase the robustness towards external disturbances proactively, Ferrolho et al.  [20, 21] incorporate a robustness metric into the trajectory optimization to compute offline plans for interactions under worst-case scenario. While these approaches are effective for end-effector control of legged mobile manipulators, they rely on an available system model and pre-specified contact schedule. These assumptions make model-based approaches vulnerable to large unmodeled disturbances (for instance, a heavy grasped object), slippages, and unknown terrain models.
@@ II-B Learning-based whole-body control
More recently, RL has shown to be a powerful alternative to model-based approaches for legged robots, providing robust control policies for locomotion on unstructured terrains [1, 3] and whole-body manipulation [9, 12, 22]. In the context of task-space tracking of a legged mobile manipulator’s arm, existing works formulate the tracking problem using different rewards and representations for the end-effector commands.
@@
Fu et al.  [9] train a combined RL policy for locomotion and manipulation. For the arm, they sample position commands in spherical coordinates and orientation commands uniformly in $\mathbb{SO}(3)$. However, their results are shown only for 3D position commands and on flat terrain. A follow-up comparison in [26] reports a poor pose tracking accuracy with this work as $22.2$\mathrm{cm}$$ and $66.22\degree$. Considering the importance of regulating forces during interactions, Portela et al.  [14] train an RL controller for tracking end-effector 3D position and 3D force simultaneously. However, they also demonstrate their work only on flat terrains. More recently, Ha et al.  [26] use a 3D Cartesian and 6D rotation representation [27] for the end-effector pose and train a low-level RL policy to track a trajectory of pose commands. While they achieve a position and orientation error of $2.12$\mathrm{cm}$$ and $3.35\degree$, respectively, the learned controller is mainly demonstrated over reference end-effector motions in front of the robot, which does not necessitate active base (leg) usage.
@@ III METHOD
We train a policy to track end-effector target poses with minimal foot displacement, intended for use alongside a locomotion policy. Figure 3 illustrates the overall training process. We use Isaac Lab [28] as a simulation environment to train our policy and deploy our controller on ALMA [4], which integrates the Anymal D robot from ANYbotics [29] with the Dynaarm from Duatic [30]. The robot has an inertial measurement unit providing body orientation and 18 actuators with joint position encoders. The main control loop and the state estimator are executed at 400 Hz while our policy runs at 50Hz on an onboard computer. While we show results on ALMA, our method remains applicable across different robot embodiments.
@@ III-B Command sampling
The distribution of the pre-sampled commands is not uniform. We fit five concentric cylinders aligned along the base z-axis. The first cylinder is solid, and each subsequent cylinder has a larger radius with a hollow section. We determine their radii by first calculating the maximum radius in the $x$-$y$ plane from the 3D positions of all pre-sampled poses and dividing it into five equal parts. The distribution of poses across these cylinders, starting from the inner one, are as follows: 51%, 21%, 13%, 6%, and 2%. To enforce more spatially uniform sampling, we randomly select one of the bins and then sample a pose from within that selected bin.
@@ III-D Action and Observation Space
The observation, $o^{t}$, is augmented with the end-effector pose command for the policy input. This command is defined as the positional difference between the current and the target keypoints of the end-effector in the base frame $\prescript{b}{}{\!p_{ee}^{\text{cmd}}}\in\mathbb{R}^{9}$, where each of the three keypoints provides a 3D position vector, resulting in a 9-dimensional vector in total. The observation does not include explicit terrain information. Nevertheless, as the robot poses are sampled from collision-free configurations (III-G) and the robot remains mostly stationary, it can deduce terrain properties indirectly.
@@ III-F Terrains and Curriculum training
The robots are trained in simulation on four procedurally generated terrains: flat, randomly rough, discrete obstacles, and stairs, as defined in [32]. Gradually increasing task difficulty during training has been shown to enhance learning [1, 35, 32]. We employ a terrain curriculum similar to the one proposed in [32], adapted for end-effector pose tracking. If the average position tracking error across the three commands within an episode is under $20$ $\mathrm{cm}$ when the task reward is active (during the final 2 seconds of each 4-second command), and the average orientation tracking error is under $20\degree$, the terrain difficulty increases after the next reset. The terrain difficulty decreases if the error exceeds $80$ $\mathrm{cm}$ and $120\degree$. Robots at the highest level are assigned a random level to prevent catastrophic forgetting.
@@ IV-A1 Pose representation comparison
Table I presents the average position and orientation errors on flat terrain for different added masses on the end-effector in simulation. Within the training range $[0,2.0]$ $\mathrm{kg}$, the tracking errors remain stable, with an average position error of $0.83$ $\mathrm{cm}$ and orientation error of $3.45\degree$. When the added mass exceeds the training distribution, the tracking performance degrades with errors reaching $15.33$ $\mathrm{cm}$ and $45.02\degree$ for a $4.5$ $\mathrm{kg}$ load. This highlights the controller’s robustness beyond the training range while also demonstrating its limitations when encountering significantly higher payloads.
@@ IV-A2 Comparison to model-based control
We compare the tracking performance of our whole-body RL policy with the model-based whole-body MPC controller from prior work [19] optimized for the same robot. For this comparison, we fine-tuned the original weight parameters to optimize whole-body behavior. To enable more natural movement, we set the MPC base reference weights to zero. We evaluate both controllers on flat terrain using the same 35 end-effector pose commands in the expanded workspace (Figure 4B). Both controllers perform similarly in terms of median errors, with the RL controller achieving $1.81$ $\mathrm{cm}$ / $1.73\degree$, and the MPC controller $2.17$ $\mathrm{cm}$ / $1.53\degree$. However, the mean errors are notably higher for the MPC controller, reaching $6.43$ $\mathrm{cm}$ / $6.88\degree$, while the RL controller maintains significantly lower values at $2.21$ $\mathrm{cm}$ / $2.01\degree$. This discrepancy results from the MPC controller’s inability to manage the trade-off between pose tracking and self-collision avoidance, causing the arm to get stuck near the base during transitions between distant poses – an issue that our RL controller effectively avoids.
Since the MPC controller was tuned specifically for flat terrain, we did not compare its performance with our whole-body RL policy on stairs. Unlike our RL policy, the MPC controller would require a terrain model to handle rough terrain, which was beyond the scope of our evaluation.
@@ IV-B1 Pose tracking accuracy
TABLE I: Average position $\bar{e}_{p}$ and orientation $\bar{e}_{o}$ errors for different added masses on the end-effector $m_{a}$ for 10000 end-effector pose commands on flat terrain in simulation.
$m_{a}$ [kg] [0 - 2.0] 2.5 3.0 3.5 4.0 4.5
$\bar{e}_{p}$ [cm] 0.83 1.18 1.89 4.77 10.69 15.33
$\bar{e}_{o}$ [deg] 3.45 6.99 10.87 22.54 36.31 45.02
@@ V CONCLUSION
We have presented a whole-body RL-based controller for a quadruped with an arm that can reach even the most difficult poses. Our controller achieves high accuracy also on rough terrain (e.g., on stairs), which we show in real-world experiments with ANYmal with an arm. Additionally, our contributions include providing a formulation for learning pose tracking that is superior to existing methods with poor accuracy or only considering position tracking. Our hardware experiments show an average tracking accuracy of $2.64~{}$\mathrm{cm}$$ for position and $3.64\degree$ for orientation on challenging terrain. Future work includes integrating 3D environment representations for self-learned collision avoidance, as in [37], and improving tracking performance under heavy, unmodeled payloads by incorporating a memory network (e.g., LSTM [38]) or a concurrent state estimation architecture [39].